# Experimental protocol for observing single quantum many-body scars with transmon qubits

### v1

@@ Abstract
Quantum many-body scars are energy eigenstates which fail to reproduce thermal expectation values of local observables in systems, where the rest of the many-body spectrum fulfils eigenstate thermalization. Experimental observation of quantum many-body scars has so far been limited to models with multiple scar states. Here we propose protocols to observe single scars in architectures of fixed-frequency, fixed-coupling superconducting qubits. We first adapt known models possessing the desired features into a form particularly suited for the experimental platform. We develop protocols for the implementation of these models, through trotterized sequences of two-qubit cross-resonance interactions, and verify the existence of the approximate scar state in the stroboscopic effective Hamiltonian. Since a single scar cannot be detected from coherent revivals in the dynamics, differently from towers of scar states, we propose and numerically investigate alternative and experimentally-accessible signatures. These include the dynamical response of the scar to local state deformations, to controlled noise, and to the resolution of the Lie-Suzuki-Trotter digitization.
@@ 1 Introduction
The discovery of quantum many-body scars (QMBSs) has represented an important twist in the theory of thermalization of isolated quantum systems, with potential applications envisioned in areas like quantum metrology [1, 2] and quantum information processing [3]. QMBSs are energy eigenstates that do not give rise to thermal expectation values in otherwise generic quantum many-body systems with almost all other energy eigenstates obeying eigenstate thermalization [4, 5, 6]. There, they weakly violate the eigenstate thermalization hypothesis [7, 8]. They are typically highly excited energy eigenstates characterized by low bipartite entanglement entropy compared to other states at similar energies.
Many scarred models feature a tower of scar states which have equidistant energies and can lead to revivals in the time evolution of certain initial states. Towers of scar states are, however, not necessarily present in all scarred models. Building on the tensor-network framework [9], we can construct models either with a tower of scar states or with a single scar. It has been shown that a scar-state phase transitions can take place in such a model with a single QMBS [10]. This phenomenon exhibits both similarities with standard ground-state phase transitions, despite taking place in a highly excited state, and novel features. QMBSs have been observed experimentally in systems with towers of scar states through the revivals of certain initial states [11, 12, 13]. However, so far there has not been an experimental observation of a single QMBS.
@@
While towers of scar states can be diagnosed from the coherent revivals in the dynamics arising from their interference, this method cannot be used for a single scar, whose experimental observation thus requires novel approaches. We propose different signatures which take advantage of the accurate single-site control and measurement available in the circuit QED toolbox. In particular, we combine controlled trotterized dynamics with the possibility to measure certain local observables as well as (Rényi) entanglement entropies. Moreover, we analyse the impact of (controlled) noise and digitization errors as a probe of the instability of the scar to thermalization.
@@ 2 Parent Hamiltonians
In Ref. [10], a family of 1D spin-1/2 models has been proposed which features a single zero-energy scar state |$\ket{S}$, based on a matrix-product-state construction. The Hamiltonian is built as the linear combination of $L$ operators ${h}_{i}$,
@@
with arbitrary coefficients $k_{i}\neq 0$. Each operator ${h}_{i}$ annihilates the scar state, ${h}_{i}\ket{S}=0$, and acts locally on three consecutive spin sites labelled by $i$, $i+1$, and $i+2$. The general expression of ${h}_{i}$ depends on two free parameters, $g\in[-1,1]$ and $a\in[0,1]$, and is given in Appendix A. In this work, we consider a fixed value $a=1$ and two different values of $g$, namely $g=\pm 1$. These choices yield a particularly simple form of ${h}_{i}$, and thus of $H$, making the model more amenable for experimental implementation. The resulting operators read as
${h}_{i}^{+}$ $=-\sigma_{i}^{z}+\sigma_{i}^{z}\sigma_{i+1}^{x},\quad$ (for $(\text{for }g=+1),$ (2)
${h}_{I}^{-}$ $=-\sigma_{i}^{z}-\sigma_{i+1}^{x}\sigma_{i+2}^{z},\quad$ (for $(\text{for }g=-1),$ (3)
@@
|$\ket{S_{+}}=\prod_{j}\ket{+}_{j},$ (4)
where $\ket{+}$ is the $+1$ eigenstate of $\sigma^{x}$. For $g=-1$, it is the 1D cluster state [14], defined as
|$\ket{S_{-}}=\left(\prod_{i=1}^{N-1}\mathrm{CZ}_{i,i+1}\right)\ket{S_{+}},$ (5)
@@ 3.1 Building blocks
Case of the $x$-polarized state— To obtain the fundamental two-qubit building block for the parent Hamiltonian (2) of the $x$-polarized state, the $i$th qubit in the chain is driven at the frequency of the $i+1$st qubit. Then, the cross-resonance Hamiltonian (6) only reproduces the addend in Eq. (2), if the coefficients in Eq. (7) fulfil the following conditions,
@@
Since $[\sigma_{i}^{z},H_{i,j}]=0$, this can be achieved by applying a (real or virtual [22]) single-qubit $z$ rotation, which does not otherwise affect the system. To meet the second condition (8b), a correction to the $c_{i,i+1}^{x}\sigma_{i+1}^{x}$ term is needed. This can be achieved by applying a resonant drive to the $i+1$st qubit [23], which generates an effective term $d_{i+1}^{x}\sigma_{i+1}^{x}$ with
@@
Case of the cluster state—The cross resonance effect can similarly be used to generate an effective three-site parent Hamiltonian for the cluster state. In this case the $i+1$st qubit is driven at the frequency of the $i$th qubit, generating the Hamiltonian
@@
with the coefficients given in Eqs. (7). Once again, additional single-qubit corrections are needed to reproduce the cluster-state parent Hamiltonian Eq. (3). A resonant drive can be applied to the $i+1$st qubit to generate the term $d_{i}^{x}\sigma_{i}^{x}$. Equivalently to Eqs. (10) and (11), the compensation condition $d_{i}^{x}=-c_{i+1,i}^{x}$ imposes
@@ 3.2 Trotterization and stroboscopic Hamiltonian
where $\omega_{1,2,3}$ denote the three transmon frequencies which are repeated every three sites. For the $x$-polarized state the order of the transmon frequencies is $[\omega_{1},\omega_{2},\omega_{3},\dots]$ and for the cluster state the order is $[\omega_{3},\omega_{2},\omega_{1},\dots]$. The order of the frequencies is chosen to obtain approximately homogeneous values of the coefficients given in Eqs. (7) throughout the spin chain, while using the same driving amplitude $\Omega$ on all sites. Spatially homogeneous values rule out potential disorder-induced effects preventing the thermalization of non-scar eigenstates. For different choices of frequency patterns, spatially homogeneous coefficients can be obtained by making $\Omega$ site-dependent and adapting it accordingly. For the Trotter step $T$ and the system size $L$, we use
@@
Here and in the following, we assume that the $x$-polarized and the cluster states can be initially prepared as the starting point of the scar-detection protocol. Assuming that all transmons are initially cooled down to their ground state, the $x$-polarized state |$\ket{x}$ is easily prepared with half-a-Rabi oscillation through a resonant microwave drive on each qubit. Starting from this state, the cluster state is obtained according to Eq. (5), by applying a controlled-$Z$ gate on each pair of neighbouring qubits. The necessary CZ gates may be performed in parallel on non-overlapping pairs and can be implemented, for fixed-frequency transmons, through the cross-resonance effect or other recently proposed native gates [35]. 1D cluster states of variable size have been prepared both with fixed-coupling superconducting qubits [24, 36] and in architectures with tuneable couplers [37].
@@ 4.2 Sensitivity to protocol distortion
where $r$ is the maximal strength of the error, $u_{i,j}^{\alpha}$ are uniformly distributed random numbers in the range $-1\leq u_{i,j}^{\alpha}\leq 1$, and $\alpha$ denotes $\alpha\in\{z,x,zx,zz\}$. This type of error does not necessarily mimic actual experimental errors in detail, but serves as a guideline for how robust the observation of the scar states is to errors in general. Similar protocol distortions may be deliberately induced by altering the noise-free value of the driving amplitude or frequency of the cross-resonance and resonant drives.
@@ 4.3 Sensitivity to coarse Trotterization
However, a different behaviour can be expected for the scar state. Indeed, recall that the scarred model is obtained by constructing operators $h_{i}$ which all annihilate the scar, Eq. (1). Meanwhile, the Lie-Suzuki-Trotter decompositions proposed in Sec. 3.2 involves stepwise Hamiltonians $\mathcal{H}_{n}$ which are sums of operators approximating $h_{i}$ on disjoint subsystems [see Eqs. (18), (12) and (17)]. It follows that the higher-order terms in $T$ in the Floquet Hamiltonian (20), given by the Baker-Campbell-Hausdorff expansion, approximately involve nested commutators of the $h_{i}$ of the form $[h_{i},\dots,[h_{j},h_{k}]\dots]]$. Since the $h_{i}$ annihilate the scar state, so do also all such nested commutators. Therefore, one can expect that, even at large $T$, the expectation values for the scar state at stroboscopic times should be weakly affected by the coarseness of the Trotterization.
@@ Acknowledgements
This work has been supported by Carlsbergfondet under Grant No. CF20-0658 and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Research Unit FOR 2414 - project number No. 277974659.
@@ Appendix A General form of the annihilation operators of the single-scar model
$|\psi_{1}\rangle\propto$ $-g|{{\uparrow}}{{\uparrow}}{{\uparrow}}\rangle-|{{\uparrow}}{{\uparrow}}{{\downarrow}}\rangle+|{{\uparrow}}{{\downarrow}}{{\uparrow}}\rangle+|{{\uparrow}}{{\downarrow}}{{\downarrow}}\rangle,$ (A.2a)
$|\psi_{2}\rangle\propto$ $-|{{\downarrow}}{{\uparrow}}{{\uparrow}}\rangle-|{{\downarrow}}{{\uparrow}}{{\downarrow}}\rangle+|{{\downarrow}}{{\downarrow}}{{\uparrow}}\rangle+g|{{\downarrow}}{{\downarrow}}{{\downarrow}}\rangle,$ (A.2b)
@@
where $a$ is a free parameter chosen between zero and one. We indicated with $\ket{{\uparrow}}$ and $\ket{{\downarrow}}$ the $+1$ and $-1$ eigenstates of $\sigma_{z}$, respectively. In this work we choose $a=1$ and $g=\pm 1$, which yields a simple form of $h_{i}$ and thus of $H$.

### v2

@@ Abstract
Quantum many-body scars are energy eigenstates which fail to reproduce thermal expectation values of local observables, in systems where the rest of the many-body spectrum fulfils eigenstate thermalization. Experimental observation of quantum many-body scars has so far been limited to models with multiple scar states evenly spaced in energy. It is thus an interesting question whether even single isolated scars, which theoretically embody the weakest possibile violation of eigenstate thermalization and may be thought to have no detectable impact in experiments, can leave a trace in measurable quantities. Moreover, single scars offer an interesting scenario for exploring the connection between quantum many-body scars and the original notion of scarring in quantum dynamical systems theory. Here we propose protocols to observe single scars in architectures of fixed-frequency, fixed-coupling superconducting qubits. We first adapt known models possessing the desired features into a form particularly suited for the experimental platform. We develop protocols for the implementation of these models, through trotterized sequences of two-qubit cross-resonance interactions, and verify the existence of the approximate scar state in the stroboscopic effective Hamiltonian. Since a single scar cannot be detected from coherent revivals in the dynamics, differently from towers of scar states, we propose and numerically investigate alternative and experimentally-accessible signatures. These include the dynamical response of the scar to local state deformations, to controlled noise, and to the resolution of the Lie-Suzuki-Trotter digitization.
@@ 1 Introduction
The discovery of quantum many-body scars (QMBSs) has represented an important twist in the theory of thermalization of isolated quantum systems, with potential applications envisioned in areas like quantum metrology [1, 2] and quantum information processing [3].
Generic quantum many-body systems are believed to fulfill the eigenstate thermalization hypothesis (ETH) [4, 5], according to which any eigenstate with finite energy density should yield thermal expectation values of local observables. This behaviour would thus reconcile the reversibility of isolated quantum dynamics with the predictions of equilibrium statistical mechanics. Known exceptions are integrable and many-body localized [6, 7, 8, 9] systems, with the latter being still under debate [10, 9]: these systems strongly violate the ETH, in the sense that even all their eigenstates could avoid thermalization, due to the presence (or emergence, in the case of many-body localization) of an extensive number of conservation laws. As an intermediate scenario, weak forms of ETH-violation have been identified [11, 12, 13], where only a few eigenstates violate the ETH, in a number which is exponentially smaller than the Hilbert-space size. This is the case of QMBSs, which are thus excited eigenstates that do not give rise to thermal expectation values in otherwise generic quantum many-body systems. QMBSs are also related to the more general phenomenon of Hilbert-space fragmentation [14], where the Hilbert space dynamically decomposes into exponentially-many disconnected subspaces with the system size, which are not associated with trivial local symmetries.
Different theoretical mechanisms have been identified, which lead to the emergence of quantum many-body scarring and which explain the rich landscape of QMBSs that has been discovered. The first phenomenology of QMBSs encountered is as a set of low-entropy eigenstates with equal energy spacing, dubbed towers of states. Towers of QMBSs have been understood to appear, for instance, in models possessing (potentially approximate) so-called spectrum generating algebras [15], where the tower can be constructed by the repeated action of a ladder-type eigenoperator of the Hamiltonian on a suitable eigenstate. The tower of states can in this case be interpreted as the creation of quasiparticles which are non-interacting (due to the equal energy spacing) on top of a given state [13, 16]. A key feature of towers of QMBSs, which enabled their observation in experiments [17, 18, 19], is that they can produce recurrences in the dynamics from certain simple initial states, e.g., in the probability of return to the initial state (Loschimdt echo). This is a consequence of the homogeneous level spacing across the tower, which makes possible for the amplitudes of the initial state on the states of the tower to interfere constructively at certain times.
Besides models with emergent towers of states, models with an arbitrary number of QMBSs can also be designed by “manually inserting” scars at the heart of the spectrum. This can be achieved, for example, by means of so-called projector-embedding methods [20, 21, 22], often combined with matrix-product-state (MPS) formulations [23]. This approach highlights that the phenomenon of quantum many-body scarring is not necessarily associated with towers of states, but isolated QMBSs can also exist. Examples include models where individual scars are annihilated by a set of quasilocal operators [20, 22], similarly to the ground states of frustration-free models, as well as frustrated [24, 25] and topologically ordered  [26, 27, 28] systems.
While towers of scar states can be diagnosed from the coherent recurrences in the dynamics, as discussed above, individual QMBSs cannot give rise to such an effect and have so far not been observed. The possibility to observe a single scar is, however, an interesting challenge for several reasons. First of all, QMBSs are by definition hard to detect, being excited states immersed in a thermal crowd, whose existence often relies on fine-tuned conditions. A natural question is, then, how weak can the violation of ETH be for actually representing an experimentally relevant and detectable phenomenon: can the ETH be violated even at the single-state level? It has further been shown that individual QMBSs can exhibit novel phenomena, such as “scar-state phase transitions” [22], where a QMBS embedded in the middle of the spectrum undergoes a phase transition while remaining a QMBS as a parameter is changed.
Finally, at an even more fundamental level, single QMBSs could be seen as closer counterparts to the original idea of quantum scarring in quantum dynamical systems [29, 30], as opposed to towers of scars. Indeed, from the perspective of quantum chaos theory, the presence of towers of equally-spaced and localized states and the related dynamical recurrences may emerge also from the quantization of locally stable orbits of the corresponding classical system associated with invariant Kolmogorov-Arnold-Moser tori [31]. It is an interesting observation that for a single quantum many-body scar this scenario can be excluded. The remarkable feature of quantum scarring in its original formulation is, actually, that a weak form of eigenstate localization is possible even in the absence of locally-stable orbits [29]. However, such a behavior is generally difficult to prove in quantum many-body systems, which usually do not have a well defined classical analog. Thus, the absence of a tower of states, which might be related simply also to analogs of locally stable KAM-tori, might provide an indication of behavior that is closer to the original idea of quantum scarring.
@@
As discussed above, the experimental observation of single scars cannot rely on coherent recurrences, differently from towers of states, and thus requires novel approaches. We propose different signatures which take advantage of the accurate single-site control and measurement available in the circuit QED toolbox. In particular, we combine controlled trotterized dynamics with the possibility to measure certain local observables as well as (Rényi) entanglement entropies. Moreover, we analyse the impact of (controlled) noise and digitization errors as a probe of the instability of the scar to thermalization.
@@ 2 Parent Hamiltonians
[…]
@@
[…]
${h}_{i}^{+}$ $=-\sigma_{i}^{z}+\sigma_{i}^{z}\sigma_{i+1}^{x},\quad$ $(\text{for }g=+1),$ (2)
${h}_{i}^{-}$ $=-\sigma_{i}^{z}-\sigma_{i+1}^{x}\sigma_{i+2}^{z},\quad$ $(\text{for }g=-1),$ (3)
@@
[…]
@@ 3.1 Building blocks
Case of the $x$-polarized state— To obtain the fundamental two-qubit building block for the parent Hamiltonian (2) of the $x$-polarized state, the $i$th qubit in the chain is driven at the frequency of the i+1i+1st qubit. Then, the cross-resonance Hamiltonian (6) only reproduces the addend in Eq. (2), if the coefficients in Eq. (7) fulfil the following conditions,
@@
Since $\sigma_{i}^{z},H_{i,j}]=0$, this can be achieved by applying a (real or virtual [40]) single-qubit $z$ rotation, which does not otherwise affect the system. To meet the second condition (8b), a correction to the $c_{i,i+1}^{x}\sigma_{i+1}^{x}$ term is needed. This can be achieved by applying a resonant drive to the i+1i+1st qubit [41], which generates an effective term $d_{i+1}^{x}\sigma_{i+1}^{x}$ with
@@
Case of the cluster state—The cross resonance effect can similarly be used to generate an effective three-site parent Hamiltonian for the cluster state. In this case the i+1i+1st qubit is driven at the frequency of the $i$th qubit, generating the Hamiltonian
@@
with the coefficients given in Eqs. (7). Once again, additional single-qubit corrections are needed to reproduce the cluster-state parent Hamiltonian Eq. (3). A resonant drive can be applied to the i+1i+1st qubit to generate the term $d_{i}^{x}\sigma_{i}^{x}$. Equivalently to Eqs. (10) and (11), the compensation condition $d_{i}^{x}=-c_{i+1,i}^{x}$ imposes
@@ 3.2 Trotterization and stroboscopic Hamiltonian
where $\omega_{1,2,3}$ denote the three transmon frequencies which are repeated every three sites. For the $x$-polarized state the order of the transmon frequencies is $\omega_{1},\omega_{2},\omega_{3},\dots]$ and for the cluster state the order is $\omega_{3},\omega_{2},\omega_{1},\dots]$. The order of the frequencies is chosen to obtain approximately homogeneous values of the coefficients given in Eqs. (7) throughout the spin chain, while using the same driving amplitude $\Omega$ on all sites. Spatially homogeneous values rule out potential disorder-induced effects preventing the thermalization of non-scar eigenstates. For different choices of frequency patterns, spatially homogeneous coefficients can be obtained by making $\Omega$ site-dependent and adapting it accordingly. For the Trotter step $T$ and the system size $L$, we use
@@
[…]
@@ 4.2 Sensitivity to protocol distortion
where $r$ is the maximal strength of the error, $u_{i,j}^{\alpha}$ are uniformly distributed random numbers in the range $1\leq u_{i,j}^{\alpha}\leq 1$, and $\alpha$ denotes $\alpha\in\{z,x,zx,zz\}$. This type of error does not necessarily mimic actual experimental errors in detail, but serves as a guideline for how robust the observation of the scar states is to errors in general. Similar protocol distortions may be deliberately induced by altering the noise-free value of the driving amplitude or frequency of the cross-resonance and resonant drives.
@@ 4.3 Sensitivity to coarse Trotterization
However, a different behaviour can be expected for the scar state. Indeed, recall that the scarred model is obtained by constructing operators $h_{i}$ which all annihilate the scar, Eq. (1). Meanwhile, the Lie-Suzuki-Trotter decompositions proposed in Sec. 3.2 involves stepwise Hamiltonians $\mathcal{H}_{n}$ which are sums of operators approximating $h_{i}$ on disjoint subsystems [see Eqs. (18), (12) and (17)]. It follows that the higher-order terms in $T$ in the Floquet Hamiltonian (20), given by the Baker-Campbell-Hausdorff expansion, approximately involve nested commutators of the $h_{i}$ of the form $h_{i},\dots,[h_{j},h_{k}]\dots]]$. Since the $h_{i}$ annihilate the scar state, so do also all such nested commutators. Therefore, one can expect that, even at large $T$, the expectation values for the scar state at stroboscopic times should be weakly affected by the coarseness of the Trotterization.
@@ Acknowledgements
This work has been supported by Carlsbergfondet under Grant No. CF20-0658 and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the Research Units FOR 2414 (projectNo. 277974659) and FOR 5688 (project No. 521530974). F. P. acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Emmy Noether Programme – project number 555842149. The authors thank the anonymous referee 2 for pointing out the connection between single quantum many-body scars and Hellers’ original concept of quantum scars as well as Roland Ketzmerick for helpful comments on this issue.
Code availability
The code used for the numerical simulations is available from the authors upon reasonable request.
@@ Appendix A General form of the annihilation operators of the single-scar model
$|\psi_{1}\rangle\propto$ $|,\displaystyle-g|{{\uparrow}}{{\uparrow}}{{\uparrow}}\rangle-|{{\uparrow}}{{\uparrow}}$$\downarrow$$\rangle$${{\uparrow}}{{\downarrow}}{{\uparrow}}\rangle+|{{\uparrow}}$$\downarrow$$\downarrow$$\rangle$, (A.2a)
[…]
@@
[…]